OpenStudy (anonymous):

Find the interval on which f(x)= 1+x^(3/2)-(3/2)x - 1 is increasing

hero (hero):

You're going to need f'(x) and f''(x) for this one

OpenStudy (anonymous):

how?:/

hero (hero):

\[f(x) = 1 + x^{3/2} - \frac {3} {2} x -1\] is that the function?

OpenStudy (anonymous):

hey hero u used to be heromilies ????

hero (hero):

If it is, then: \[f'(x) = \frac {3} {2}x^{1/2} - \frac {3} {2}\] \[f''(x) = \frac {3} {4x^{1/2}}\]

hero (hero):

Yes, I used to be heromiles

OpenStudy (anonymous):

oh i was wondering who u wer wen everybody was talkin bout u

OpenStudy (anonymous):

I think this is correct but everybody take a look and let me know. Go to this link at scribblar.com http://www.scribblar.com/kuub1xb

OpenStudy (anonymous):

it wont load on my computer

OpenStudy (anonymous):

mmm...oh well it would take me forever to type all that. and to show the graph...I got wher -1<x<1....is that correct?

OpenStudy (anonymous):

yeah i think so it looks right to me (:(:

hero (hero):

Rhino, your post takes me to a sign-in page

OpenStudy (anonymous):

just make up a name and sign in...no password needed

OpenStudy (anonymous):

sorry i couldnt help u more /:

OpenStudy (anonymous):

bye guys(:(:(:(:

hero (hero):

Wow, we medaled each other simultaneously